OpenStudy (anonymous):

there is a linear relationship between the number of people in a group and the cost to enter a museum. the museum charges $20 for two people and 28 for three. what is the equation that relates the number of people in a group to the cost of entering the museum? (slope-intercept form)

OpenStudy (anonymous):

Two points (2,20) (3,28) y-y1=m(x-x1) 28-20=m(3-2) 8=m y-20=8(x-2) y=8x-16+20 y=8x+4 Check: x=2 y=8*2+4=20 x=3 y=8*3+4=28